A rectangular area of 16 meters by 12 meters, is surrounded by a road 3 meters wide, the area of the road is?
Answer by Theo(13342)

the length of the rectangular area plus the road on each end is 16 + 6 = 22 meters.
the width of the rectangular area plus the road on each end is 12 + 6 = 18 meters.

the area of the rectangular area plus the width of the road is 22 * 18 = 396 square meters.

the area of the rectangular area by itself is 16 * 12 = 192 square meters.

the area of the road is the total area minus the rectangular area which is equal to 396 - 192 = 204 square meters.

the width of the road adds 3 meters to the length of the rectangular area on both sides which increases the total length by 6 meters.

the width of the road adds 3 meters to the width of the rectangular area on both sides which increases the total width by 6 meters.

the total area minus the inside area is equal to the surrounding area which is the area of the road.